OpenStudy (dls):

The distance of the line 2x-3y=4 from the point(1,1) in the direction of the line x+y=1 is

OpenStudy (kunal):

solve the two equations and get their intersection point..... use distance formula to find the req distance

OpenStudy (dls):

they are not intersecting

OpenStudy (anonymous):

Solve 2x-3y=4 and x+y=1 Find x and y ........ After Use Distance Formula

OpenStudy (anonymous):

as wat kunal said

OpenStudy (kunal):

no their intersection point is 7/5,-2/5

OpenStudy (kunal):

(-2,2) is not the intersection point

OpenStudy (dls):

now im geting \[\LARGE \sqrt{\frac{53}{25}}\]

OpenStudy (dls):

\[\LARGE \sqrt{ ( \frac{-2}{5}-1 )^2+(\frac{7}{5}-1)^2}\]

OpenStudy (kunal):

yes it is....

OpenStudy (dls):

Its wrong sorry

OpenStudy (kunal):

then whats the answer??

OpenStudy (dls):

\[\LARGE \sqrt{2}\]

OpenStudy (dls):

they dont even intersect lol what are u guys saying..we cant solve them together

OpenStudy (dls):

they both are parallel lines..and parallel lines NEVER intersect

OpenStudy (sirm3d):

is this from analytic geometry?

OpenStudy (dls):

co-ordinate

OpenStudy (sirm3d):

do you know the formula for the distance of a point from a line?

OpenStudy (kunal):

see... the equations are 2x-3y=4 and x+y=1 their slope are not the same therefore they got to intersect

OpenStudy (dls):

i know the one for perpendicular one..

OpenStudy (sirm3d):

okay. we'll use that. first, what is the slope of the given line \(2x-3y=4\)?

OpenStudy (dls):

2/3

OpenStudy (sirm3d):

right. so the line \(\bot\) has slope __ ?

OpenStudy (dls):

-3/2

OpenStudy (sirm3d):

good. now you have the slope of the \(\bot\) line and a point on that line, \((1,1)\). Can you find the equation of that line?

OpenStudy (dls):

\[\large (y-1)=\frac{-2}{3}(x-1)\]

OpenStudy (dls):

3y+2x-5=0

OpenStudy (sirm3d):

close. it should be \[(y-1)=\frac{-3}{2}(x-1)\]or\[3x+2y-5=0\]

OpenStudy (dls):

ah sorry.!

OpenStudy (sirm3d):

now we have two intersecting lines, \(2x-3y=4\) and \(3x+2y-5=0\). can you find the point of intersection of the two lines?

OpenStudy (dls):

y=-22/14 or \[\LARGE y= \frac{-\pi}{2}\] :P

OpenStudy (dls):

where pi=22/7

OpenStudy (anonymous):

@DLS see first of all find the eq of line parallel to x+y=1 and passing through (1,1) the line passing through 1,1 and parallel to x+y=1 is x+y=2 now we have to find the point of intersection of 2x-3y=4 and x+y=2 which is (2,0) now we need to find the distance between (1,1) and (2,0) using distance formula thus reqd distance =sqt(2) ans is sqrt(2) units ...

OpenStudy (anonymous):

@DLs first check is the ans sqrt(2)

OpenStudy (dls):

yes

OpenStudy (dls):

I had a hint on this question..which can be possibly shorter..

OpenStudy (sirm3d):

ooh, i didn't get that part, "in the direction of the line".

OpenStudy (dls):

Any point on the line through(1,1) and parallel to x+y=1 is \[p(1+ rcos \theta,1+r \sin \theta) where \tan \theta=-1.\] Use the fact that P is on the line \[2x-3y=4\] and find |r|.

OpenStudy (dls):

what does it mean :/

OpenStudy (anonymous):

@DLS then follow the method as i have done ... i am certain abt it

OpenStudy (shubhamsrg):

the question is that there is a line parallel to x+y=1, passing through (1,1) , and the line intersects 2x-3y =4 somewhere . we need to find the distance between (1,1) and that common co-ordinate..

OpenStudy (shubhamsrg):

sorry i didnt notice @matricked has already explained that part..

OpenStudy (dls):

why did we find the point of intersection? how do we know it will intersect

OpenStudy (shubhamsrg):

since the lines are not parallel, ofcorse they'll intersect.. !

OpenStudy (dls):

i see..thanks! everyone